### v1

@@ Abstract
With the rapid advancement of Quantum Machine Learning (QML), the critical need to enhance security measures against adversarial attacks and protect QML models becomes increasingly evident. In this work, we outline the connection between quantum noise channels and differential privacy (DP), by constructing a family of noise channels which are inherently $\epsilon$-DP: $(\alpha,\gamma)$-channels. Through this approach, we successfully replicate the $\epsilon$-DP bounds observed for depolarizing and random rotation channels, thereby affirming the broad generality of our framework. Additionally, we use a semi-definite program to construct an optimally robust channel. In a small-scale experimental evaluation, we demonstrate the benefits of using our optimal noise channel over depolarizing noise, particularly in enhancing adversarial accuracy. Moreover, we assess how the variables $\alpha$ and $\gamma$ affect the certifiable robustness and investigate how different encoding methods impact the classifier’s robustness.
@@ I Introduction & Related Work
Quantum noise channels, such as depolarizing and phase damping noise in NISQ devices, are used as sources of stochastic noise. This noise helps achieve DP by exploiting the natural error processes of these devices [4, 15, 33, 27, 10]. With this argumentation, Weber et al. [27] have provided a relationship among quantum hypothesis testing and adversarial robustness. In addition, Hirche et al. [15] has provided a relationship between quantum DP and the quantum hockestick divergence. Angrisani et al. [4] is possibly one of the most noteworthy recent publications; they have provided a more general framework for the robustness upper bound of quantum noise channels by defining the neighbourhood concept in terms of the Schatten-norm and providing a more general upper bound for noise channels in terms of a depolarizing channel as well as a single qubit Pauli channel. This broader framework was not included in our work as we initially focused on defining such a family of noise channels. The linear nature of quantum channels and their relationship to semi-definite programming has been already considered, e.g. Guan et al. [14], to attempt to construct an optimal noise channel that achieves epsilon-DP using semi-definite programming. Nonetheless, the extent to which these techniques can defend classifiers against practical adversarial input manipulations remains uncertain.
As indicated in our earlier work [29], the effectiveness of these approaches in providing robustness varies with the factor of depolarization noise. This raises the question which noise channel is the most suitable for a given classification problem. On a pathway to answer this question, we develop a family of noise channels, so called $(\alpha,\gamma)$-channels, which describe noise channels naturally facilitating DP. In addition, we offer theoretical derivations for the bounds of depolarizing noise [33] and random rotations [16] from our general framework of ($\alpha$, $\gamma$)-channels. This indicates that a broader set of channels can be described in our framework. Next, we add a construction of an optimal quantum channel as a positive semi-definite constrained optimization problem. This construction allows the experimental evaluation of a depolarizing channel against its optimally constructed counterpart, providing a more definite statement about the possible utility of quantum noise channels for DP and adversarial robustness against evasion attacks.
In summary, the contributions of our work are:
  • The introduction of ($\alpha$, $\gamma$)-channels as a family of noise channels which suffice $\epsilon$-DP, along with formal bound derivations for depolarizing noise and random rotations.
  • Construction of an optimal ($\alpha$, $\gamma$)-channel trough a semi definite program. The maximum robustness ratio is obtained for a quantum classifier and a given sample.
  • Experimental evaluation of depolarizing noise channels compared to optimal noise channels and evaluation between amplitude and angle encoding on three datasets.
@@ II-A Quantum Adversarial Robustness
Definition 1 (Adversarial Robustness).
Consider a classifier $f$ that maps an input $x$ to an output $y$. The adversarial robustness at input $x$ can be defined as the smallest amount of perturbation $\delta$ needed to change the model’s prediction:
$\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x+\delta)\neq\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x),$
@@
In the context of QC, a study by Lu et al. [21] found that Quantum Neural Networks (QNN) can fall prey to these targeted perturbations. As part of their work, the quantum version of the Fast Gradient Sign Method (FGSM) [13] was introduced. Thus, let us denote $x\in\mathbb{R}^{n}$ as (classical) features, $\mathcal{L}$ as loss function incorporating the QNN weights and $\nabla_{x}$ as gradient. In the context of FGSM a adversarial budget of $\epsilon$ is assumed, which then allows to obtain an adversarial example as follows:
$x_{adv}=\Pi_{x+\mathcal{S}}(x+\epsilon\cdot\mathrm{sign}(\nabla_{x}\mathcal{L})),$
where $\Pi_{x+\mathcal{S}}(\cdot)$ clips the value towards the space $\mathcal{S}$ of adversarial perturbations around $x$. Typically, the space of adversarial perturbations is defined by a norm around the input $x$. The fact that QNNs are vulnerable to adversarial perturbations motivates the discussion about specific measures to counteract adversarial examples. In this work, we consider robustness accuracy as the network’s ability to main stable predictions against adversarial attacks.
Definition 2 (Robustness Accuracy).
Let $\mathcal{D}=\{x_{i},{y}_{i}\}_{i=1,\dots,N}$ be a dataset and $a(\cdot,f,\epsilon):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be an adversarial model that tries to find an adversary within the perturbation budget $\epsilon$. Then, the robustness accuracy of the classifier $f$ on $\mathcal{D}$ under attack $a(\cdot,f,\epsilon)$ is defined as:
$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}({y}_{i}=f(a(x_{i},f,\epsilon)))$ (1)
Where $\mathbbm{1}$ denotes the indicator function which corresponds to 1 if the argument is true, else is returning 0.
II-B Quantum DP
DP, with parameters $(\epsilon,\delta)$, enhances data analysis and machine learning protection against adversaries. Lower $\epsilon$ values increase privacy by limiting what an algorithm reveals about individual data points, while $\delta$ represents the small chance that this privacy might fail. Together, $(\epsilon,\delta)$-DP makes it hard for adversaries to exploit data, safeguarding data integrity and privacy against sophisticated attacks. In the following, we focus entirely on $\epsilon$-DP, which implies that $\delta=0$ throughout this work.  [^1] This is justified, since typically $\epsilon$-DP is used for classification problems. Nevertheless as pointed out by Angrisani et al. [4] $\epsilon$-DP cannot describe depolarizing noise entirely.
Formally for QC, two neighbouring Hilbert states $\rho$ and $\sigma$ are assumed, which are bound by a distance measure. Typically, the trace distance:
@@ II-B Quantum DP
A quantum algorithm $A$ (for instance, a classifier) is assumed to be $\epsilon$-DP if, for a given set of outcomes S of $A(x)$, the following equation holds:
$\text{Pr}(A(\rho)\in S)\leq e^{\epsilon}\text{Pr}(A(\sigma)\in S).$ (3)
Since computations on quantum devices are dictated by a CPTP map (a quantum channel $\mathcal{E}$) and outcomes can be characterized by a set of Positive Operator-Valued Measurements (POVMs) $\{\Pi_{k}\}$ [^2] Set of orthogonal projectors., as alternative, one can abstract the effect of a quantum classifier as follows:
$\text{Pr}(A(\sigma)\in S)=y_{k}(\sigma)=\text{Tr}\left[\Pi_{k}\mathcal{E}(\sigma)\right].$ (4)
This reformulation leads to the definition of quantum $\epsilon$-DP as:
@@ II-C1 Bound on Depolarizing Noise
A depolarizing noise channel is characterized by the parameter $p$, defined as follows:
@@ II-C2 Bound on random rotations
where $t=(h+1)\cdot 2^{(-n)}$ is the noise level. Their bound is of particular interest, since it is independent of the underlying device.
@@ II-D Choi representation of noise channels
$\mathcal{E}(\rho)=\sum_{k}K_{k}\rho K_{k}^{\dagger}\quad\text{where}\quad\sum_{k}K_{k}K_{k}^{\dagger}=I.$ (9)
@@
where |$\ket{\phi^{+}}$ represents the fully entangled state and $A$ the auxiliary system which we assume to always have the same dimensionality as our $n$-qubit system. On a further note, we refer to the Choi matrix that describes the identity channel as identity Choi matrix: $\mathcal{J}_{\mathcal{I}}=\ket{\phi^{+}}\bra{\phi^{+}}$ |.
Next, it is necessary to ensure that $\mathcal{J}_{\mathcal{E}}$ is Completly Positive (CP) and Trace Preserving (TP) for our resulting quantum channel to be Completly Positive and Trace Preserving (CPTP). This can be achieved by asserting that the Choi matrix is positive semi-definite ($\mathcal{J}_{\mathcal{E}}\succeq 0$) and that tracing out the system describing the noise channel lead to an identity matrix of the size of the subsystem ($\text{Tr}_{2}\left[\mathcal{J}_{\mathcal{E}}\right]=I_{A}$). Thus, we can construct the application of a noise channel as  [^3] See [32] for a visualization using tensor networks. [18]:
@@ III Introduction of ($\alpha$, $\gamma$)-Channels
In this section, we introduce a family of quantum noise channels, which allow us to derive $\epsilon$-DP bounds and robustness guarantees. In reviewing the proofs concerning $\epsilon$-DP conducted by Du et al. [10], Huang et al. [16] and Zhou and Ying [33], we observe that their bounds share similarities, particularly in their use of the property that trace preserving operations are contractive with respect to the trace distance, formally:
@@
In addition, they typically manage to establish bounds for the minimum probability of the noise channel, i.e:
$y_{k}(\sigma)>0.$ (15)
Essentially, their proofs rely on the fact that quantum noise channels are contractive, and some may offer a lower bound on the smallest possible POVM. This reasoning inspires the definition of a family of noise channels that can be demonstrated to meet the requirements of $\epsilon$-DP.
@@
Intuitively, any $(\alpha,\gamma)$-channel has a contraction factor of at most $\alpha$ and any input $\sigma$ result in at most $\gamma$ after a POVM.
@@ III-C Certifiable robustness on $(\alpha,\gamma)$-channels
This section builds on the robustness bound from Du et al. [10] and generalizes it to an $(\alpha,\gamma)$-channel. We focus on the infinite sampling case for a hypothetical noise model that can only be simulated, not physically implemented. Refer to Du et al. [10] for details on the finite sampling case.
@@
Next, one can simply use $\epsilon=\ln\left[1+\frac{\alpha\tau_{D}}{\gamma}\right]$, which will lead to:
@@ IV Construction of optimal ($\alpha$, $\gamma$)-Channels
We consider a single data point (later expanded to a set) represented by a density matrix $\sigma=\sigma(x)$. A pretrained quantum classifier described by a unitary $U$ acting on $n$-qubits and measured on $d$ qubits by $2^{d}$ POVM-measurements: $\{\Pi_{k}\}$. We assume a true label $C$ (w.l.o.g.). For brevity, we denote $\sigma_{D}=U\sigma U^{\dagger}$ and abbreviate this by $\sigma$. Refer to Eq. 4 for the POVM measurement definition: $y_{k}(\sigma)=\text{Tr}\left[\Pi_{k}\mathcal{E}_{(\alpha,\gamma)}(\sigma)\right]$, where $\mathcal{E}_{(\alpha,\gamma)}$ represents the noise channel with parameters $\alpha$ and $\gamma$. Given a threshold $B=\frac{y_{C}(\sigma)}{y_{k\neq C}(\sigma)}>e^{2\epsilon}$, we recognize that $\epsilon$ is fixed by the chosen noise channel. Hence, to maximize classifier robustness, a natural approach is to maximize the fraction $B$. More specifically we can formulate this as the following optimization problem:
@@ V-A Classification of Depolarizing Noise against optimal noise
(b) Iris (Angle)
(c) BC (Amplitude)
(d) PID (Amplitude)
Figure 1: Adversarial accuracy given a FGSM attack for the various datasets used in this study. The dashed line indicates the adversarial accuracy’s of the best case noise channel’s. The red dashed vertical line corresponds to the smallest $L_{\infty}$ distance between two data samples of the opposite class. Note that we used a single $(\alpha,\gamma)$-channel, since they showed the tendency to overlap (See Appendix C).
@@ V-C Comparison of amplitude and angle embedding
Figure 3: Normalized adversarial accuracy’s between amplitude and angle encoding. The adversarial attacks where done using FGSM with a $L_{\infty}$-norm.
@@ VI Conclusion
In this work, we have defined a family of noise channels, which are certifiable robust. We have hereby related robustness of QNNs to the contraction of the algorithm, to the smallest possible label assignment as well as to the encoding strategy used by the quantum classifier. Our SDP formulation, while theoretically useful for constructing optimal noise channels (maximizing the fraction $B$), suffers from numerical scaling limitations in practical applications. Nevertheless, the study reveals several interesting implications and future research directions.
Firstly, the framework could be extended to evaluate other noise sources beyond depolarizing noise, potentially leading to a broader understanding of quantum noise and its impact on $\epsilon$-DP.
Secondly, while a small amount of depolarizing noise shows promise as a defense against evasion attacks, it remains unclear if noise inherent to NISQ devices can truly provide robustness. Our findings, along with previous work, suggest that increasing depolarization leads to decreased certifiability.
Thirdly, the concept of noise channels as a general model for quantum computation, coupled with the existence of certifiable robust channels, opens exciting avenues. Constructing a variational ansatz with inherently certifiable behavior would be a significant advance. Such an ansatz could be easier to optimize compared to finding the optimal noise model (a convex optimization problem), despite the computational challenges arising from exponentially growing Hilbert spaces. This could pave the way for theoretically grounded adversarial robustness in quantum classifiers. Additionally, the observed improvement in robustness due to encoding suggests that quantum classifiers might be inherently robust when data point separation is effective. This combination of factors could lead to a paradigm shift, where quantum classifiers are recognized for their inherent robustness that scales with data separability.

### v2

@@ Abstract
Quantum Machine Learning (QML) is rapidly evolving. Recent studies suggest that quantum noise could improve classifier robustness. To verify this claim, we developed a semidefinite programming model that, given a dataset and a QML ansatz, constructs the most robust noise channel possible. In our small-scale experiments, we first explore the range of behaviors generated by this optimized noise channel. Secondly, we show that, despite recent claims, the direct impact of noise is modest compared to the substantial robustness improvements achieved through increasing the number of qubits. Aside from that, we utilize our framework to assess different noise channels in terms of their robustness and certifiability.
@@ I Introduction & Related Work
Quantum noise channels, such as depolarizing and phase damping noise in NISQ devices, are used as sources of stochastic noise. This noise helps achieve DP by exploiting the natural error processes of these devices [4, 15, 33, 27, 10]. With this argumentation, Weber et al. [27] have provided a relationship among quantum hypothesis testing and adversarial robustness. In addition, Hirche et al. [15] has provided a relationship between quantum DP and the quantum hockestick divergence. Angrisani et al. [4] is possibly one of the most noteworthy recent publications; they have provided a more general framework for the robustness upper bound of quantum noise channels by defining the neighbourhood concept in terms of the Schatten-norm and providing a more general upper bound for noise channels in terms of a depolarizing channel as well as a single qubit Pauli channel. This broader framework was not included in our work as we initially focused on defining such a family of noise channels. The linear nature of quantum channels and their relationship to semidefinite programming has been already considered, e.g. Guan et al. [14] used semi-definite programming to derrive the optimal bound of a quantum classifier and detect adversarial examples. Nonetheless, the extent to which these techniques shield quantum classifiers from practical adversarial input manipulations remains unclear.
As indicated in our earlier work [29], the effectiveness in terms of adversarial accuracy of these approaches in providing robustness varies with the factor of depolarization noise. To make this more clear, a quantum classifier with depolarization noise strength of one is inherently robust, but is equivalent to flipping a coin. This circumstance allows asking for the best depolarization noise strenght or more general the noise channel that is as robust as possible. To answer the question of the most robust noise channel, we develop a family of noise channels, so called $(\alpha,\gamma)$-channels. We found that every channel in this group is naturally robust because of its origin. Next, to show the effectiveness of our strategy, we offer theoretical derivations for the bounds on robustness for depolarizing noise [33] and random rotations [16] from our general framework of ($\alpha$, $\gamma$)-channels. This indicates that a broader set of channels can be described in our framework. As a final step, we add a construction of an optimal quantum channel as a positive semidefinite constrained optimization problem. This construction allows the experimental evaluation of a depolarizing channel against its optimally constructed counterpart, providing a more definite statement about the possible utility of quantum noise channels for DP and adversarial robustness against evasion attacks.
@@ II-A Quantum Adversarial Robustness
$\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x+\delta)\neq\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x),$ (1)
@@
In the context of QC, a study by Lu et al. [21] found that Quantum Neural Networks (QNN) can fall prey to these targeted perturbations. As part of their work, the quantum version of the Fast Gradient Sign Method (FGSM) [13] was introduced. On a high level FGSM computes the gradient of the loss with respect to the input and then slightly alters the input in the direction that most increases the loss, effectively creating an adversarial example by using a small pertubation.
The fact that QNNs are vulnerable to those small perturbations motivates the discussion about specific measures to counteract adversarial examples. To measure the effectiveness of approaches, we consider robustness accuracy as the network’s ability to main stable predictions against adversarial attacks. We define this robustness accuracy as the fraction of samples that where correctly predicted by the network under attack.
II-B Quantum Differential Privacy
[…]
$\text{Pr}(M(D)\in S)\leq e^{\epsilon}\text{Pr}(M(D^{\prime})\in S),$ (2)
where Pr indicates probability. Before introducing the concept of quantum differential privacy, we want to elaborate on the concept of distance and the concept of processing in quantum computing. Concerning the distance, it can be defined as the trace distance between two density matrices $\rho$ and $\sigma$:
@@ II-B Quantum Differential Privacy
Computations on quantum devices are dictated by a quantum channel $\mathcal{E}$, which is a Completly Positive and Trace Preserving (CPTP) map. Outcomes can be characterized by a set of Positive Operator-Valued Measurements (POVMs) $\{\Pi_{k}\}$. Hence, we can express the outcome of a quantum classifer, which we assume aligning with the definition above to be a mechanism:
$\text{Pr}(M(\sigma)\in S)=y_{k}(\sigma)=\text{Tr}\left[\Pi_{k}\mathcal{E}(\sigma)\right]$ (4)
This leads to the definition of quantum $\epsilon$-DP on a noisy quantum algorithm $M$, by stating that for every pair of density matrices ($\sigma,\rho$) within a specific distance $\tau(\sigma,\rho)$, it holds for all possible measurement outcomes, that:
@@ II-C1 Bound on Depolarizing Noise
A depolarizing noise channel [^2] Note that this defines a global noise channel. is characterized by the parameter $p$, defined as follows:
@@ II-C2 Bound on random rotations
where $t=\frac{h+1}{\sqrt[n]{2}}$ is the noise level.  [^3] Note that to the best of our knowledge $h$ was not formally defined in Huang et al. [16]. We concluded that it originated as $h$ from the proof in [16, Appendix: Proof of Lemma 1]. The lower bound was given as $\frac{(h+1)^{n}}{2}$, which we equated to $t^{n}=\frac{(h+1)^{n}}{2}$ in the definition of their $\tau_{D}$ and solved accordingly. Their bound is of particular interest, since it is independent of the underlying device.
@@ II-D Choi representation of noise channels
$\mathcal{E}(\rho)=\sum_{i}K_{i}\rho K_{i}^{\dagger}\quad\text{where}\quad\sum_{i}K_{i}K_{i}^{\dagger}=I.$ (9)
@@
where $\otimes$ is a tensor product and |$\ket{\phi^{+}}$ represents the fully entangled state and $A$ the auxiliary system which we assume to always have the same dimensionality as our $n$-qubit system. On a further note, we refer to the Choi matrix that describes the identity channel as identity Choi matrix: $\mathcal{J}_{\mathcal{I}}=\ket{\phi^{+}}\bra{\phi^{+}}$ |.
Next, it is necessary to ensure that $\mathcal{J}_{\mathcal{E}}$ is CPTP for our resulting quantum channel. This can be achieved by asserting that the Choi matrix is positive semidefinite ($\mathcal{J}_{\mathcal{E}}\succeq 0$) and that tracing out the system describing the noise channel lead to an identity matrix of the size of the subsystem ($\text{Tr}_{2}\left[\mathcal{J}_{\mathcal{E}}\right]=I_{A}$). Thus, we can construct the application of a noise channel as  [^4] See Wood et al. [32] for a visualization using tensor networks. [18]:
@@ III Introduction of ($\alpha$, $\gamma$)-Channels
In this section, we introduce a family of quantum noise channels, which allow us to derive $\epsilon$-DP bounds and robustness guarantees. In reviewing the proofs concerning $\epsilon$-DP conducted by Du et al. [10], Huang et al. [16] and Zhou and Ying [33], we observe that their bounds share similarities, particularly in their use of the property that trace preserving operations are contractive with respect to the trace distance, more formally:
@@
In addition, they typically manage to establish bounds for the minimum probability of the noise channel, i.e: $y_{k}(\sigma)>0$. This reasoning inspires the definition of a family of noise channels that can be demonstrated to meet the requirements of $\epsilon$-DP.
@@
Intuitively, any $(\alpha,\gamma)$-channel has a contraction factor of at most $\alpha$. Any input state $\sigma$ will result in at most $\gamma$ when measured. Note that each measurement is defined by their POVM Projector $\Pi_{k}$.
@@ III-C Certifiable robustness on $(\alpha,\gamma)$-channels
This section builds on the robustness bound from Du et al. [10] and generalizes it to an $(\alpha,\gamma)$-channel. We focus on the infinite sampling case since we base our assumptions on a hypothetical model, which can be simulated but not implemented.
@@
Next, one can simply use $\epsilon=\ln\left[1+\frac{\alpha\tau_{D}}{\gamma}\right]$, which will lead to a certifiable distance defined as:
@@ IV Construction of optimal ($\alpha$, $\gamma$)-Channels
We consider a single data point (later expanded to a set) represented by a density matrix $\sigma=\sigma(x)$. A pretrained quantum classifier described by a unitary $U$ acting on $n$-qubits and measured on $d$ qubits by $2^{d}$ POVM-measurements: $\{\Pi_{k}\}$. We assume a true label $C$ (w.l.o.g.). For brevity, we denote $\sigma_{D}=U\sigma U^{\dagger}$ and abbreviate this by $\sigma$. Each class propability is defined by the POVM-Measurement: $y_{k}(\sigma)=\text{Tr}\left[\Pi_{k}\mathcal{E}_{(\alpha,\gamma)}(\sigma)\right]$, where $\mathcal{E}_{(\alpha,\gamma)}$ represents the noise channel with parameters $\alpha$ and $\gamma$. Given a threshold $B=\frac{y_{C}(\sigma)}{y_{k\neq C}(\sigma)}>e^{2\epsilon}$, we recognize that $\epsilon$ is fixed by the chosen noise channel. Hence, to maximize classifier robustness, a natural approach is to maximize the fraction $B$. More specifically we can formulate this as the following optimization problem:
@@ V-A Classification of Depolarizing Noise against optimal noise
(b) BC (Amplitude)
(c) PID (Amplitude)
Figure 1: The plot shows adversarial accuracy under an FGSM attack across all evaluated datasets: the blue dashed curve, $E_{(\alpha,\gamma)}$, represents adversarial accuracy with our best-case noise channel, the red dashed vertical line indicates the smallest $L_{\infty}$ distance between any two samples from opposite classes, and the colored curves $E_{dep}$ depict adversarial accuracy when a depolarization noise channel is applied at increasing noise levels. We include only one $(\alpha,\gamma)$-channel because the individual channels largely overlap (See Appendix C).
@@ V-C Comparison of amplitude and angle embedding
Figure 3: Normalized adversarial accuracy’s between amplitude and angle encoding. The adversarial attacks where done using FGSM with a $L_{\infty}$-norm. The blue dashed curve, $E_{(\alpha,\gamma)}$, represents adversarial accuracy with our best-case noise channel, and the colored curves $E_{dep}$ indicate adversarial accuracy when a depolarization noise channel is applied at increasing noise levels.
@@ VI Conclusion
In this work, we have defined a family of noise channels, which are certifiable robust. We have hereby related robustness of QNNs to the contraction of the algorithm, to the smallest possible label assignment as well as to the encoding strategy used by the quantum classifier. Our SDP formulation, while theoretically useful for constructing optimal noise channels, suffers from numerical scaling limitations in practical applications.
Our framework could be extended to evaluate other noise sources beyond depolarizing noise, potentially leading to a broader understanding of quantum noise and its impact on $\epsilon$-DP. Next, while a small amount of depolarizing noise shows promise as a defense against evasion attacks, it remains unclear if noise inherent to NISQ devices can truly provide robustness. Our findings, along with previous work, suggest that increasing depolarization leads to decreased certifiability.
The concept of noise channels as a general model for quantum computation, coupled with the existence of certifiable robust channels, opens exciting avenues. Constructing a variational ansatz with inherently certifiable behavior would be a significant advance. Such an ansatz could be easier to optimize compared to finding the optimal noise model (a convex optimization problem), despite the computational challenges arising from exponentially growing Hilbert spaces. This could pave the way for theoretically grounded adversarial robustness in quantum classifiers.
At last we want to emphasize that based on this study optimizing for a robust noise channel yields a minor advantage compared to the larger increase of the robustness budget when scaling the amount of qubits.